# ProGeny II: the impact of libraries and model configurations on inferred galaxy properties in SED fitting

### v1

@@ 5.3 Impact of the SPL Initial Mass Function
Making further use of the flexibility of ProGeny, we conduct a number of tests to isolate the impact of different Initial Mass Functions (IMFs) [^13] The code to produce the ProGeny SPLs using each of the implemented IMFs is available online at https://rpubs.com/asgr/1235271. on the SED fitting output. We note that while Robotham & Bellstedt (in prep) made comparisons against a Salpeter (1955) IMF, we do not include this in our comparisons, as it is generally regarded to be too bottom heavy for low-$z$ galaxy SED fitting. For consistency, all IMFs are constructed using lower and upper mass limits of 0.1 $\rm{M}_{\odot}$ and 100 $\rm{M}_{\odot}$ respectively.
@@
Fig. 22 shows the derived CSFH of the fiducial run using ProGeny with a Chabrier (2003) IMF, as well as the equivalent when assuming a Kroupa (2002) IMF. It is notable that the variation between these two is exceptionally small - smaller even than the impact of assuming the metallicity to be constant or evolving. The main discernable difference between a Chabrier and Kroupa IMF in terms of the recovered CSFH, is the slightly higher inferred SFRs in the last $\sim$ 2 Gyr when a Kroupa IMF is assumed (although this difference is well within the uncertainties of the observed CSFH). The corresponding impact on the derived stellar masses is shown in the top panel of Fig. 23. Because the Kroupa IMF has slightly more low-mass stars than the Chbrier IMF, it is unsurprising that the stellar msses derived using the Kroupa IMF are systematically higher by around 0.03 dex. Similarly, we find that the estimated SFR values are also higher using Kroupa by around 0.03 dex.
@@
To assess the impact of a more “exotic” treatment of the IMF, we also produce a simple stellar population using ProGeny that evolves in time according to the Lacey et al. (2016) prescription. In this prescription, a simple temporal evolution is provided through a star burst dominated top-heavy IMF at high redshift (used by Lacey et al. 2016 for the “starburst” mode), transitioning to a Kennicutt (1983) IMF at a lookback time of 10 Gyr (used by Lacey et al. 2016 for the “quiescent” mode). Both the starburst and quiescent mode IMF prescriptions are broadly similar to a Kroupa IMF, with differences at the low-mass end toward a more bottom-heavy IMF for the starburst mode, and more bottom-light for the quiescent mode (see Robotham & Bellstedt in prep for more details about this implementation).
@@
The impact of this temporally evolving IMF is seen to be greater than simply the swap from a Chabrier to a Kroupa IMF (particularly in the last 5 Gyr, where the evolving IMF is seen to systematically recover lower SFRs). Of note as well, is that while the transition of IMFs at a lookback time of 10 Gyr in the Lacey prescription is somewhat arbitrary and abrupt (selected to coincide with cosmic noon), there is no corresponding impact of this seen in the CSFH, in terms of a discontinuity. This is however the transition point between two epochs that demonstrate different behaviour. In the older epoch Lacey-derived CSFH has a greater SFRD than the Chabrier-derived one, while in the younger epoch the Lacey CSFH has a lower SFRD than the Chabrier one. The impact on the derived stellar masses is seen to be substantial, in the bottom panel of Fig. 23, with a systematic increase in 0.158 dex. As is evident from the lower SFRD at present dy in Fig. 22, the corresponding SFRs are also recovered to be systematically 0.134 dex lower than with the assumption of a Chabrier IMF.
@@ 6.1 CSFH impact from SPL sub-ingredients
The second panel shows that other modifications result in a SFRD at the present day more similar to the value from D’Silva et al. (2023), however all except the Lacey IMF modification produce a value consistent within observational uncertainty. The SFRD at the CSFH peak (third panel) is best recovered by the fiducial ProGeny implementation, as is the age of the peak (shown in the fourth panel). The total amount of mass formed in each modification relative to the D’Silva et al. (2023) CSFH is then shown in the bottom panel.
Figure 24: Equivalent of Fig. 17 to characterise the impact on the derived CSFH when modifying only sub-ingredients of the ProGeny SPL. The orange points represent the fiducial implementation of ProGeny using the MIST isochrones, the C3K stellar atmospheres, and a Chabrier IMF (identical to the black ProGeny point plotted in Fig. 17). Each other point point to a result after a single modification in the ProGeny SPL.
@@ 6.2 Ranking of modelling assumptions
If retrieving accurate stellar masses is the primary aim of SED fitting, then the greatest bias will result from fixing the modelled metallicity to solar. While the impact of changing IMF from Chabrier to Kroupa is relatively small, the potential impact from assuming a more exotic IMF evolution is high (0.16 dex). SPL selection is the next most impactful choice, with a change to BPASS causing the greatest offset of $-0.15$ dex. The change of SFH parametrisation to b5 SFH creates a maximum median offset of 0.11 dex (consistent with similar estimates recorded in the literature). The specific modifications within an SPL (specifically the changing of stellar atmospheres and isochrones) have lower impacts, with maximum offsets of -0.045 and -0.018 respectively. Finally, the change to modelling a metallicity history as constant versus evolving has a negligible impact on stellar masses.
@@
If instead retrieving the SFRs is the primary aim of SED fitting, then we show that again the greatest bias is caused by fixing the modelled metallicity to solar. The bias is systematically positive, with the smallest offset identified when using the CB19 SPL, and the greatest offset found with M05 (followed closely by BPASS). While the impact of a Kroupa versus Chabrier IMF on the SFRs is small, an exotic form of IMF evolution results in the second-largest impact on SFRs after the fixed $Z_{\odot}$ assumption. These assumptions are followed by the b5 SFH parametrisation. While an offset of up to 0.11 dex can be recovered through use of b5 SFH, the large MADs indicate that the scatter caused by the SFH parametrisation is also large (also evident from Fig. 8). SPL choice has a noticeable impact on SFRs, and then the impacts of stellar atmospheres, isochrones, and constant Z are much smaller.
@@
While the half-mass age does not capture the complexities of star formation histories, Table 4 also presents the ranked list of assumptions that impact the recovered ages. This ranked list differs slightly from the other global parameters. Once again, the fixed metallicity assumption has the largest impact, underestimating ages by up to 0.47 dex. The impact of b5 SFH is the second largest, with an up to 0.12 dex overestimate recorded for BPASS, CB19 and ProGeny. The jump down in biases is smaller for the SPL, which can cause a variation of 0.1 dex in age values. Stellar atmospheres and isochrones cuase a similarly small offset, followed by the constant Z assumption, and finally the variation in IMF can be seen to have a minor impact on the recovered half-mass age.
@@
While Table 4 ranks the factors that impact the recovery of global galaxy parameters, we rank the factors separately that impact the ability to recover the CSFH in Table 5. It is notable that this ranked list is different, with the constant Z assumption driving one of the largest deviations from the CSFH (which Fig. 17 shows is caused in combination with the M05 SPL), and the impact of stellar atmospheres and isochrones within the SPL creating a similar impact to that of the b5 SFH. By far the smallest impact on the CSFH recovery is caused stellarby the IMF selection.
@@
Table 5: The relative difference range in CSFH Mean Absolute Deviations (MAD), when compared against the CSFH of D’Silva et al. (2023). Modifications are ranked from greatest to least median impact on the derived CSFH.
@@
$Z_{\odot}$ [$0.16$, $0.34$] 0.21 1
Constant Z [$0.10$, $0.37$] 0.18 2
b5 SFH [$0.16$, $0.21$] 0.18 3
Atmosphere [$0.10$, $0.21$] 0.17 4
SPL [$0.10$, $0.17$] 0.13 5
Isochrone [$0.10$, $0.21$] 0.11 6
IMF [$0.09$, $0.14$] 0.10 7
@@ 6.3 Modelling assumptions not considered in this work
Another aspect in SED fitting that will impact the ability to recover realistic properties, is the inclusion of burstiness in a SFH. Recent work has highlighted the potential impact of burstiness (for example Iyer et al., 2019, 2024; Sun et al., 2023; Narayanan et al., 2024; Haskell et al., 2024), however the difficulty in fitting burstiness for individual galaxies means that for now this is limited to studying averaged behaviour over populations.
@@
Finally, this study has focussed solely on a sample of low-redshift galaxies. When SED-fitting techniques are applied to high-$z$ galaxies, then the total amount of cosmic time requiring modelling is substantially reduced, thereby changing the nature of the systematics at play. High-redshift fitting may therefore present a much more favourable application for stepwise parametric SFHs, where individual bins are not required to cover such large time periods Indeed, much recent SFH fitting of high-$z$ galaxies has been conducted in this way, for example Nanayakkara et al. (2024), however functional parametric forms are still proving popular, for example Carnall et al. (2023). Furthermore, the nature of star formation is much more extreme in the high-redshift Universe, and it is quite likely that dust properties vary (see for example Shivaei et al., 2024).
@@ 7 Conclusions
We have conducted an SED-fitting analysis using the code ProSpect on a volume-complete sample of 7,862 galaxies from the GAMA survey, to characterise the impact of SPL selection against the impact of SED-fitting modelling choices like the metallicity implementation and SFH parametrisation. Furthermore, we quantify the impact of sub-SPL changes, including isochrone selection atmospher selection, and the chosen IMF implementation. A number of conclusions have been reached from this analysis:
  • The greatest bias we find in estimating galaxy properties from SED fitting is caused by fixing the galaxy metallicity to $Z_{\odot}$, which has the potential to underestimate stellar masses by 0.27 dex, overestimate SFRs by 0.2 dex, and underestimate half-mass ages by 0.47 dex.
  • The biases introduced through the choice of stellar poulation library (SPL) in SED fitting are of a similar magnitude to those introduced by choices like the SFH parametrisation, in terms of the derived properties of stellar mass, SFR, and half-mass age. In other words, the scatter between SPLs with the same physics (in terms of IMF, isochrones, stellar atmospheres) is generally larger than most modifications of the physics itself.
  • The evolving metallicity prescription is found to have relatively little impact on global galaxy properties like stellar mass and SFR, however it removes systematic biases in the recovered cosmic star formation history (CSFH) and produces more realistic uncertainties in the recovered SFHs. To recover the observed CSFH, it is essential to include an evolving metallicity.
@@
  • The isochrone selection within each SPL is a major factor influencing the differences in recovered SFHs at the SED fitting level, whereas the choice of stellar atmospheres produces a much smaller shift in the SFHs.
@@ Acknowledgements
The authors acknowledge the contributions made by Luke Davies and Simon Driver through various discussions. ASGR and SB acknowledge support from the ARC Future Fellowship scheme (FT200100375). SB acknowledges funding by the Australian Research Council (ARC) Laureate Fellowship scheme (FL220100191). This work was supported by resources provided by the Pawsey Supercomputing Centre with funding from the Australian Government and the Government of Western Australia. In total, around 174,000 CPU hours were required for the testing and final computation of all outputs presented in this work.

### v3

@@ 5.3 Impact of the SPL Initial Mass Function
Figure 25: The Initial Mass Functions (IMFs) tested in this work (top panel), and their residual relative to the fiducial Chabrier IMF (bottom panel). All IMFs have been normalised to show the same amount of stellar mass.
@@
Making further use of the flexibility of ProGeny, we conduct a number of tests to isolate the impact of different Initial Mass Functions (IMFs) [^16] The code to produce the ProGeny SPLs using each of the implemented IMFs is available online at https://rpubs.com/asgr/1235271. on the SED fitting output. We note that while Robotham & Bellstedt (2024) made comparisons against a Salpeter (1955) IMF, we do not include this in our comparisons, as it is generally regarded to be too bottom heavy for low-$z$ galaxy SED fitting. For consistency, all IMFs are constructed using lower and upper mass limits of 0.1 $\rm{M}_{\odot}$ and 100 $\rm{M}_{\odot}$ respectively. We plot all IMFs in Fig. 25, where the IMF itself is shown in the top panel, and the residual to the fiducial Chabrier IMF in the bottom panel. In the bottom panel we highlight the four regions of parameter space where the terms ‘top-heavy’, ‘top-light’, ‘bottom-heavy’, and ‘bottom-light’ are used to describe the shape of th IMF relative to Chabrier.
@@
Fig. 26 shows the derived CSFH of the fiducial run using ProGeny with a Chabrier (2003) IMF, as well as the equivalent when assuming a Kroupa (2002) IMF. It is notable that the variation between these two is exceptionally small - smaller even than the impact of assuming the metallicity to be constant or evolving. The main discernable difference between a Chabrier and Kroupa IMF in terms of the recovered CSFH, is the slightly higher inferred SFRs in the last $\sim$ 2 Gyr when a Kroupa IMF is assumed (although this difference is well within the uncertainties of the observed CSFH). The corresponding impact on the derived stellar masses is shown in the top panel of Fig. 27. Because the Kroupa IMF has slightly more low-mass stars than the Chabrier IMF, it is unsurprising that the stellar msses derived using the Kroupa IMF are systematically higher by around 0.03 dex. Similarly, we find that the estimated SFR values are also higher using Kroupa by around 0.03 dex.
@@
To assess the impact of a more “exotic” treatment of the IMF, we also produce a simple stellar population using ProGeny that evolves in time according to the Lacey et al. (2016) prescription. In this prescription, a simple temporal evolution is provided through a star burst dominated top-heavy IMF at high redshift (used by Lacey et al. 2016 for the “starburst” mode, presented as the dashed grey line in Fig. 25), transitioning to a Kennicutt (1983) IMF at a lookback time of 10 Gyr (used by Lacey et al. 2016 for the “quiescent” mode, presented s the dashed black line in Fig. 25). Note from Fig. 25 that while most top-heavy IMFs are bottom-light (and vice versa), the starburst mode from Lacey et al. (2016) occupies both the top-heavy region and the bottom-heavy, after transitioning through the bottom-light quadrant. Both the starburst and quiescent mode IMF prescriptions are more top-heavy than the Kroupa IMF, with differences at the low-mass end toward a more bottom-heavy IMF for the starburst mode and more bottom-light for the quiescent mode.
@@
The impact of this temporally evolving IMF is seen to be greater than simply the swap from a Chabrier to a Kroupa IMF (particularly in the last 5 Gyr, where the evolving IMF is seen to systematically recover lower SFRs). Of note as well, is that while the transition of IMFs at a lookback time of 10 Gyr in the Lacey prescription is somewhat arbitrary and abrupt (selected to coincide with cosmic noon), there is no corresponding impact of this seen in the CSFH, in terms of a discontinuity. This is however the transition point between two epochs that demonstrate different behaviour. In the older epoch Lacey-derived CSFH has a greater SFRD than the Chabrier-derived one, while in the younger epoch the Lacey CSFH has a lower SFRD than the Chabrier one. The impact on the derived stellar masses is seen to be substantial, in the bottom panel of Fig. 27, with a systematic increase in 0.158 dex. As is evident from the lower SFRD at present day in Fig. 26, the corresponding SFRs are also recovered to be systematically 0.134 dex lower than with the assumption of a Chabrier IMF.
@@ 6.1 CSFH impact from SPL sub-ingredients
The second panel shows that other modifications result in a SFRD at the present day more similar to the value from D’Silva et al. (2023), however all except the Lacey IMF modification produce a value consistent within observational uncertainty. The SFRD at the CSFH peak (third panel) and the age of the peak (shown in the fourth panel) is marginally better recovered by the Kroupa and evolving Kroupa modifications than the fiducial implementation. The total amount of mass formed in each modification relative to the D’Silva et al. (2023) CSFH is then shown in the bottom panel.
Figure 28: Equivalent of Fig. 20 to characterise the impact on the derived CSFH when modifying only sub-ingredients of the ProGeny SPL. The orange points represent the fiducial implementation of ProGeny using the MIST isochrones, the C3K stellar spectra, and a Chabrier IMF (identical to the black ProGeny point plotted in Fig. 20). Each other point point to a result after a single modification in the ProGeny SPL.
@@ 6.2 Ranking of modelling assumptions
If retrieving accurate stellar masses is the primary aim of SED fitting, then the greatest bias will result from fixing the modelled metallicity to solar. While the impact of changing IMF from Chabrier to Kroupa is relatively small, the potential impact from assuming a more exotic IMF evolution is high (0.16 dex). SPL selection is the next most impactful choice, with a change to BPASS causing the greatest offset of $-0.15$ dex. The change of SFH parametrisation to b5 SFH creates a maximum median offset of 0.11 dex (consistent with similar estimates recorded in the literature). The specific modifications within an SPL (specifically the changing of stellar spectra and isochrones) have lower impacts, with maximum offsets of -0.045 and -0.018 respectively. Finally, the change to modelling a metallicity history as constant versus evolving has a negligible impact on stellar masses.
@@
If instead retrieving the SFRs is the primary aim of SED fitting, then we show that again the greatest bias is caused by fixing the modelled metallicity to solar. The bias is systematically positive, with the smallest offset identified when using the CB19 SPL, and the greatest offset found with M05 (followed closely by BPASS). While the impact of a Kroupa versus Chabrier IMF on the SFRs is small, an exotic form of IMF evolution results in the second-largest impact on SFRs after the fixed $Z_{\odot}$ assumption. These assumptions are followed by the b5 SFH parametrisation. While an offset of up to 0.11 dex can be recovered through use of b5 SFH, the large MADs indicate that the scatter caused by the SFH parametrisation is also large (also evident from Fig. 10). SPL choice has a noticeable impact on SFRs, and then the impacts of stellar spectra, isochrones, and constant Z are much smaller.
@@
While the half-mass age does not capture the complexities of star formation histories, Table 4 also presents the ranked list of assumptions that impact the recovered ages. This ranked list differs slightly from the other global parameters. Once again, the fixed metallicity assumption has the largest impact, underestimating ages by up to 0.47 dex. The impact of b5 SFH is the second largest, with an up to 0.12 dex overestimate recorded for BPASS, CB19 and ProGeny. The jump down in biases is smaller for the SPL, which can cause a variation of 0.1 dex in age values. Stellar spectra and isochrones cause a similarly small offset, followed by the constant Z assumption, and finally the variation in IMF can be seen to have a minor impact on the recovered half-mass age.
@@
While Table 4 ranks the factors that impact the recovery of global galaxy parameters, we rank the factors separately that impact the ability to recover the CSFH in Table 5. In contrast to the MAD values presented in Fig. 20, this table demonstrates the deviation between the base CSFH and the modified version, as a mechanism of conveying relative impact. It is notable that this ranked list is different, with the constant Z assumption driving one of the largest deviations from the CSFH (where M05 is most discrepant), and the median impact of stellar spectra and isochrones within the SPL creating a similar impact to that of the constant Z modification, followed closely by the b5 SFH. By far the smallest impact on the CSFH recovery is caused by the IMF selection.
@@
Table 5: The relative difference range in CSFH Mean Absolute Deviations (MAD), when compared against the fiducial CSFH. Modifications are ranked from greatest to least median impact on the derived CSFH.
@@
$Z_{\odot}$ [$0.11$, $0.24$] 0.17 1
Constant Z [$0.07$, $0.29$] 0.10 2
Stellar Spectra [$0.07$, $0.12$] 0.10 3
Isochrone [$0.07$, $0.12$] 0.10 3
b5 SFH [$0.09$, $0.25$] 0.09 5
SPL [$0.06$, $0.13$] 0.08 6
IMF [$0.02$, $0.09$] 0.02 7
  • * Comparison is conducted by modifying the configuration for each given SPL.
@@ 6.3 Modelling assumptions not considered in this work
Another aspect in SED fitting that will impact the ability to recover realistic properties, is the inclusion of burstiness in a SFH. Recent work has highlighted the potential impact of burstiness (for example Iyer et al., 2019, 2024; Sun et al., 2023; Narayanan et al., 2025; Haskell et al., 2024), however the difficulty in fitting burstiness for individual galaxies means that for now this is limited to studying averaged behaviour over populations.
@@
Finally, this study has focussed solely on a sample of low-redshift galaxies. When SED-fitting techniques are applied to high-$z$ galaxies, then the total amount of cosmic time requiring modelling is substantially reduced, thereby changing the nature of the systematics at play. High-redshift fitting may therefore present a much more favourable application for stepwise parametric SFHs, where individual bins are not required to cover such large time periods. Indeed, much recent SFH fitting of high-$z$ galaxies has been conducted in this way, as for example in Nanayakkara et al. (2025), however functional parametric forms are still proving popular, as for example in Carnall et al. (2023). Furthermore, the nature of star formation is much more extreme in the high-redshift Universe, and it is quite likely that dust properties vary (see for example Shivaei et al., 2024).
@@ 7 Conclusions
We have conducted an SED-fitting analysis using the code ProSpect on a volume-complete sample of 7,862 galaxies from the GAMA survey, to characterise the impact of SPL selection against the impact of SED-fitting modelling choices like the metallicity implementation and SFH parametrisation. Furthermore, we quantify the impact of sub-SPL changes, including isochrone selection, stellar spectra selection, and the chosen IMF implementation. A number of conclusions have been reached from this analysis:
  • The greatest bias we find in estimating galaxy properties from SED fitting is caused by fixing the galaxy metallicity to $Z_{\odot}$, which has the potential to underestimate stellar masses by 0.27 dex, overestimate SFRs by 0.2 dex, and underestimate half-mass ages by 0.47 dex. Furthermore, this assumption typically produces the greatest impact on the inferred cosmic star formation history (CSFH).
  • The biases introduced through the choice of stellar poulation library (SPL) in SED fitting are of a similar magnitude to those introduced by choices like the SFH parametrisation, in terms of the derived properties of stellar mass, SFR, and half-mass age. While some contributing factors relate to underlying physics that changes (such as a different IMF or method of deriving stellar isochrones), other aspects of the implementation also have substantial impact that are not rooted in physics (such as the interpolation mechanisms used).
  • The evolving metallicity prescription is found to have relatively little impact on global galaxy properties like stellar mass and SFR, however it removes systematic biases in the derived CSFH and produces more realistic uncertainties in the recovered SFHs. To recover the observed CSFH, it is essential to include an evolving metallicity.
@@
  • Both the isochrone selection and the stellar spectra selection are significant factors in influencing the differences in recovered SFHs at the SED fitting level, each with a potential impact equivalent to the constant metallicity assumption and the change in SFH parametrisation.
@@ Acknowledgements
The authors thank the anonymous referee for their detailed reading of the paper, which has resulted in multiple improvements. The authors also acknowledge the contributions made by Luke Davies and Simon Driver through various discussions and the anonymous referee via their helpful recommendations. ASGR and SB acknowledge support from the ARC Future Fellowship scheme (FT200100375). SB acknowledges funding by the Australian Research Council (ARC) Laureate Fellowship scheme (FL220100191). This work was supported by resources provided by the Pawsey Supercomputing Centre with funding from the Australian Government and the Government of Western Australia. In total, around 174,000 CPU hours were required for the testing and final computation of all outputs presented in this work.